OpenStudy (anonymous):

Solve x + 4y = 6 -y + z = 2 x + z = 5

OpenStudy (anonymous):

SUbtract 3rd equation from 2nd equation

OpenStudy (anonymous):

you will get an equation in x and y as -(x+y)=-3 so,x+y=3

OpenStudy (anonymous):

solve this with first x+4y=6..................(1) x+y=3.....................(2) subtract 2 from 1:: 3y=3 y=1 put the value of y in 2 ,we get: x+1=3 x=2

OpenStudy (anonymous):

put the value of x in 3rd equation ie x+z=5 ie 2+z=5 z=3

OpenStudy (anonymous):

ya right????????